OpenStudy (anonymous):

a salesman earns commission on units sold. the more units he sells, the more commission he earns on a particular unit. for the first unit,he earns $20. for each subsequent unit, he earns 2% more.how much does the sales man earn on the 25th unit?

OpenStudy (anonymous):

\(20\times 1.02\) for the second unit \(20\times (1.02)(1.02)=20\times (1.02)^2\) for the third \(20\times (1.02^2)(1.02)=20\times (1.02)^3\) for the fourth any idea for the 25th?

OpenStudy (anonymous):

i pluged in 24 where the 2 is and pluged in 25 where the 3 is. but i get nothing. 20(1.02^24)(1.02) = 20(1.02)^25

OpenStudy (anonymous):

pattern should suggest in the 25 one you have \(20(1.02)^{24}\) need a calculator for his

OpenStudy (anonymous):

about $32.17 http://www.wolframalpha.com/input/?i=20%281.02%29^24

OpenStudy (anonymous):

oh i see, thanks